What is the concentration of a Ca(OH)2 solution if 15.0 ml of 0.500 M H3PO4 solution is required to completely neutralize 22.5 ml of the Ca(OH)2 solution?
1
Expert's answer
2013-05-01T11:47:46-0400
3Сa(OH)2 + 2H3PO4 = Ca3(PO4)2 + 6H2O
C(M) = v/V
C(M)-molar concentration of the solution;
v - molar mass, mole;
V - volume of the solution.
v = m/M
m - mass of the substance, grams;
M - molar mass of the substance.

From the task we know that 15.0 ml of 0.500 M H3PO4 solution is required to completely neutralize 22.5 ml of the Ca(OH)2.
The amount of H3PO4 could be calculated as:
v = C(M) x V
v(H3PO4) = 0.5 x 0.015 = 0.0075 mol
According to the equation the total amount of Сa(OH)2 is:
v(Сa(OH)2) = v(H3PO4)/2*3 = 0.0075/2*3 = 0.011 mol

So the concentration of Сa(OH)2 is:
C(Сa(OH)2) = v/V = 0.011/0.0225 = 0.48 M